• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Kelly

    Log entry time 00:04:10 on July15,2001

    Entry number 66194

    keyword=swing shift summary

    Summary for swing shift
    ===================
    The first two runs were taken last shift, but we need to keep a running total
    for the present configuration and include them below.
    .
    left_run.......good e-........right_run......good e-..........net.......grand total
    2023............1.78k............21023............1.54k.............3.32k.......3.32k
    2024............2.26k............21024............2.18k.............4.44k.......7.76k
    2025............9.63k............21025............9.06k...........18.69k.....26.45k
    2026............5.87k............21026............5.79k...........11.66k.....38.11k
    2027............9.43k............21027............8.86k...........18.29k.....56.40k
    2028............1.67k............21028............4.10k.............5.77k.....62.17k
    2029............8.34k............21029............7.91k...........16.25k.....78.42k
    2032............4.66k............21032............4.61k.............9.37k.....87.79k
    .
    .
    1815: epr=46.2%...........nmr=44.07%
    2030: frequent RF trips for last hour or two. MCC claims to be investigating, but
    their response was not convincing or encouraging.
    2045: no triggers in left arm; all trigger HVs off. End run 2028 and reset.
    2215: epr=46.0%..........nmr =45.1%
    2245: Piotr is playing with charge asymmetry feedback during runs 2030-2031,
    which are omitted from running total.
    2355: MCC takes beam away to adjust optics for Hall B.